# FedRTS: Federated Robust Pruning via Combinatorial Thompson Sampling

### v1

@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Neural Network Pruning
Neural network pruning (Mozer & Smolensky, 1988; LeCun et al., 1989; Janowsky, 1989) optimizes network structures by removing redundant components like connections (Han et al., 2015), neurons (Yu et al., 2018), or filters (Li et al., 2016). Traditional pruning methods focus on inference, requiring dense training followed by importance-based pruning and retraining. These importance scores are derived from weight magnitudes (Janowsky, 1989; Han et al., 2015), gradients (Mozer & Smolensky, 1988; Molchanov et al., 2019b), Fisher information (Singh & Alistarh, 2020), and other variants (Louizos et al., 2018; Yu et al., 2018). However, training a dense model first increases computational and memory costs.
Therefore, Recent works propose dynamic sparse training (Mocanu et al., 2018; Dettmers & Zettlemoyer, 2019; Evci et al., 2020) to bypass the dense training by iteratively adjusting sparse topologies during training while maintaining a fixed model size. Yet, existing methods rely on deterministic adjustments, leading to high variance and resource-intensive searches, exacerbated in federated learning under data heterogeneity and partial client availability. To address this, we propose TSAdj, which employs probabilistic adjustments for a more robust and stable model topology.
@@ A.2 Federated Pruning
Federated Pruning is a technique that applies neural network pruning within the Federated Learning (FL) process. To avoid the requirements of dense training in traditional pruning methods, some previous methods determine the sparse model topology before federated training and do not adjust it during the FL process, such as SNIP (Lee et al., 2018) and Synflow (Tanaka et al., 2020). This approach of pruning at initialization may overlook crucial data information, potentially leading to suboptimal model structures.
Dynamic sparse training methods, such as those employed by PruneFL (Jiang et al., 2022), adjust the model during FL rounds. However, these methods require the entire model gradient to be sent to the server to ensure final performance, resulting in high communication costs. FedDST (Bibikar et al., 2022) and FedSGC (Tian et al., 2024) alleviate this issue by applying dynamic sparse training on clients. FedTiny (Huang et al., 2023) and FedMef (Huang et al., 2024) require clients only to upload the Top gradients. Despite these advancements, these methods adjust the model topology based on limited information from the currently selected clients, leading to issues of greedy adjustment and unstable topology. FedRTS addresses these challenges by applying TSAdj, significantly mitigating the problems associated with limited client information and enhancing the robustness of model adjustments by leveraging probabilistic distributions.
@@ A.3 Combinatorial Multi-Arm Bandit
The Combinatorial Multi-Armed Bandit (CMAB) problem extends the classical MAB to combinatorial settings, where rewards depend on selected arm combinations. (Gai et al., 2012) introduced the CMAB framework with linear rewards and proposed a UCB-type algorithm that allows approximation oracles for decision-making. This work established the foundation for CMAB analysis.  (Chen et al., 2013, 2016) and  (Wang & Chen, 2017) generalized the CMAB framework by introducing probabilistically triggered arms and relaxing reward assumptions to monotonicity. They proposed the CUCB algorithm, achieving regret bounds with approximation oracles and enabling broader applications of CMAB.
Despite these advancements, UCB-based algorithms, such as CUCB, rely on assumptions like monotonicity for reward functions. However, these assumptions may not hold in more complex scenarios, such as non-monotonic rewards or high-dimensional combinatorial action spaces.  (Wang & Chen, 2018) pointed out that these limitations restrict the practicality of UCB algorithms, particularly in settings where reward functions are highly nonlinear or lack smoothness.
Thompson Sampling (TS) offers greater flexibility by avoiding monotonicity. TS has been shown to achieve regret bounds comparable to CUCB, as demonstrated by  (Komiyama et al., 2015) and  (Wang & Chen, 2018), while  (Perrault et al., 2020) further improved its performance coefficients.  (Kong et al., 2021) extended the theoretical understanding of CMAB by conducting a detailed hardness analysis on regret and bounds when using greedy oracles. Their study quantified the impact of greedy oracles on regret and provided analytical results for scenarios where exact solutions are unavailable. Additionally, their analysis addressed the convergence of TS in combinatorial settings, highlighting its robustness in handling complex reward structures.
@@ B.1 TSAdj
In each loop $t$, the aggregated weights $W^{agg}_{t+1}$ will be calculated in each loop when the server receives the weights $W^{n}_{t+1}$ from each client $n\in C_{t}$, and TSAdj is invoked to adjust the sparse topology $m_{t}$ only in the outer loop. The outcomes $X_{t,i}$ for link $m_{i}$ are computed as described in Equation 3, which balances global and client-specific contributions. Specifically, outcomes $X^{agg}_{t,i}$ and $X^{n}_{t,i}$ are calculated based on Equations 4 and 5 on the server, using the aggregated weights $W_{t+1}$ and the uploaded weights $W^{n}_{t+1}$, while $X^{n}_{t,i}$ will be updated in the outer loop using the indices of the top $K-\kappa$ gradients $\mathcal{I}_{t+1}^{n}$ from each client $n$ based on Equation 6. Outcomes $X^{agg}_{t,i}$ and $X^{n}_{t,i}$ are combined to compute the final outcome $X_{t,i}$ for link $m_{i}$ in Equation 3, where $\gamma\in[0,1]$ controls the trade-off between global and client-specific contributions, and $p_{n}$ reflects the relative weight of client $n$ based on the size of its local dataset $D_{n}$. A larger $\gamma$ emphasizes client-specific observations, while a smaller $\gamma$ prioritizes global contributions.
The prior probability distribution $P$ for each link $m_{i}$ in the topology $m$ is modeled as a beta distribution parameterized by $(\alpha_{i},\beta_{i})$. At initialization, $(\alpha_{i},\beta_{i})$ are set to small positive values (e.g., $\alpha_{i}=\beta_{i}=1$) to represent a uniform prior, ensuring that all links have an equal probability of activation at the outset. The beta parameters $(\alpha_{i},\beta_{i})$ for each link $m_{i}$ are then updated incorporating with the computed outcomes $X_{t,i}$, via Equation 7, where $\lambda$ controls the trade-off between exploration and exploitation by adjusting the impact of the current outcomes on the beta distribution’s update. A larger $\lambda$ increases the influence of the current loop’s outcomes $X_{t,i}$, accelerating the concentration of the Beta distribution towards a certain probability. Conversely, a smaller $\lambda$ preserves uncertainty to encourage exploration. Over time, as $\alpha_{i}$ and $\beta_{i}$ grow under the influence of $\lambda$ and $X_{t,i}$, the Beta distribution becomes sharper and more concentrated, reflecting an increasingly certain probability of selecting $m_{i}$ as an active link in the topology $m$.
@@ B.2 FedRTS
The training process is conducted over $T_{end}$ loops. At the beginning of each training loop $t$, the server employs a random sampling mechanism to select a subset of clients $C_{t}$ from the total $\mathcal{N}$ available clients for participation in the training process. This sampling process is conducted uniformly at random, ensuring that each client, indexed by $n$ from $1,\dots,\mathcal{N}$, has an equal probability of being chosen. This randomized partial participation mechanism not only avoids bias in client selection but also ensures scalability and reduces communication overhead, making it well-suited for large-scale distributed systems.
@@
  Input: Initial model $W_{0}$, sparsity $d^{\prime}$, total loops $T_{end}$, adjustment interval $\Delta T$
  Output: Final sparse model $W_{T_{end}},m_{T_{end}}$
@@
  for each loop $t$=0, 1, 2, …, $T_{end}-1$ do
@@
     Broadcast server model ($W_{t}$, $m_{t}$) to all clients $n\in C_{t}$
@@
        Receive model $W_{t}$ and mask $m_{t}$ from the server
        $W^{n}_{t+1}\leftarrow$ perform local training on $D_{n}$
@@
           Compute gradients $G^{n}_{t+1}$ respect to loss function
           $\mathcal{I}_{t+1}^{n}\leftarrow\{i|i\in\text{Top}(|G^{n}_{t+1}|,K-\kappa)\}$
@@
        $W^{agg}_{t+1},m_{t+1}=\text{TSAdj}(\{W^{n}_{t+1},\mathcal{I}_{t+1}^{n}|n\in C_{t}\}$)
@@
        $m_{t+1}\leftarrow m_{t}$
        $X^{agg}_{t}\leftarrow$ calculated by Eq.4 based on $W^{agg}_{t+1}$
        $X^{n}_{t}\leftarrow$ calculated by Eq.5 based on $W_{t+1}^{n}$
        $X_{t}\leftarrow(1-\gamma)X^{agg}_{t}+\gamma\sum_{n\in C_{t+1}}p_{n}X^{n}_{t}$
@@
After completing $T_{end}$ training loops, the server outputs the final globally aggregated weights $W_{T_{end}}$ and an optimized mask $m_{T_{end}}$, both derived from the iterative training involving all participating clients. The weights $W_{T_{end}}$ represent the learned parameters that effectively accommodate heterogeneous data distributions, while the topology $m_{T_{end}}$ preserves the robust and reliable structure of the model.
@@ C.1 Greedy Optimal Action
Theorem. (Upper Bound) Under assumptions 3.13.2 and 3.3 and with outcomes defined in Eq. 8, the regret $Reg(T)$ of TSAdj can be upper bounded by:
$\begin{split}Reg(T)&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2BK\epsilon)^{2}}\\ &+\left(2K+4\langle W\rangle+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}\\ &=O\left(\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{\Delta_{max}L^{2}\log T}{\Delta^{2}_{s,k}}\right).\end{split}$ (12)
for any $\epsilon$ such that $\forall s\not=S^{*}_{1}$ and $i\notin S^{*}_{k},\Delta_{s,k}>2BK\epsilon$, where $U$ is a universal constant.
Proof. Our proof follows (Kong et al., 2021). For any arm $s\not=s^{*}_{1}$, we apply the exploration price $F(i)$ as
@@
Therefore, at $t$-th iteration, there are two event
@@
Based on Eq. 14 and 15 , the regret $Reg(T)$ can be devided into three terms as
$\begin{split}Reg(T)&\leq\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{\neg A_{t}\land\neg B_{t}}\Delta_{S_{t}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{A_{t}}\Delta_{S_{t}}\right]\\ &+\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{B_{t}}\Delta_{S_{t}}\right].\end{split}$ (16)
We provide the Lemma 18,  24 and  27 to bound these three terms one by one. Therefore, we can obtain
$\begin{split}Reg(T)&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2BK\epsilon)^{2}}\\ &+\left(2K+4\langle W\rangle+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}\end{split}$ (17)
@@
$\begin{split}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{\neg A_{t}\land\neg B_{t}}\Delta_{S_{t}}\right]&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2BK\epsilon)^{2}}\\ &+\left(2K+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}\end{split}$ (18)
@@
$\begin{split}\mathbb{E}&\left[\sum_{t=1}^{T}\mathbf{1}_{\neg A_{t}\land\neg B_{t}}\Delta_{S_{t}}\right]\leq\sum_{k=1}^{K}\mathbb{E}\left[\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{\neg A_{t}\land\neg B_{t}\land C_{t,k}}\right]\\ &+\sum_{k=1}^{K}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{s^{*}_{k}=s_{t,k}\land||\xi_{t,s^{*}_{k}}-\mu_{s^{*}_{k}}||_{\infty}\leq\epsilon}\right]\cdot\Delta_{\max},\end{split}$ (19)
@@
$\begin{split}&\sum_{k=1}^{K}\mathbb{E}\left[\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{\neg A_{t}\land\neg B_{t}\land C_{t,k}}\right]\\ &\leq\sum_{k=1}^{K}\mathbb{E}\left[\sum_{s\notin S^{*}_{k}}\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{s=s_{t,k}\land N_{t,s}\leq F(s)}\right]+\frac{UK}{\epsilon^{6}}\Delta_{max}\\ &\leq\mathbb{E}\left[\sum_{s\not=s^{*}_{1}}\sum_{t=1}^{T}\sum_{k=1}^{K}\Delta_{S_{t}}\cdot\mathbf{1}_{s=s_{t,k}\land N_{t,s}\leq F(s)}\right]+\frac{UK}{\epsilon^{6}}\Delta_{max}\\ &\leq\mathbb{E}\left[\sum_{s\not=s^{*}_{1}}\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{s\in S_{t}\land N_{t,s}\leq F(s)}\right]+\frac{UK}{\epsilon^{6}}\Delta_{max}\\ &\leq\sum_{s\not=s^{*}_{1}}F(s)\Delta_{s}^{\max}+\frac{UK}{\epsilon^{6}}\Delta_{max}\\ &\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2BK\epsilon)^{2}}+\frac{UK}{\epsilon^{6}}\Delta_{max},\end{split}$ (21)
@@
$\begin{split}\sum_{k=1}^{K}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{s^{*}_{k}=s_{t,k}\land||\xi_{t,s^{*}_{k}}-\mu_{s^{*}_{k}}||_{\infty}\leq\epsilon}\right]&\Delta_{\max}\\ \leq 2K+\frac{8K}{\epsilon^{2}}\Delta_{\max}&.\end{split}$ (22)
@@
$\begin{split}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{\neg A_{t}\land\neg B_{t}}\Delta_{S_{t}}\right]&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2BK\epsilon)^{2}}\\ &+\left(2K+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}.\end{split}$ (23)
@@
$\begin{split}&\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{A_{t}}\Delta_{S_{t}}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{A_{t}}\right]\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{|\xi_{t,i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}}\right]\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\land|\xi_{t,i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\right)\Delta_{\max}\\ &=\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\\ &\cdot\mathbb{P}\left(|\xi_{t,i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\middle|N_{t,i}=\tau\right)\Delta_{\max}.\end{split}$ (25)
@@
$\begin{split}&\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\\ &\cdot\mathbb{P}\left(|\xi_{t,i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\middle|N_{t,i}=\tau\right)\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\cdot 2\exp(-3\log T)\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\frac{2}{T}\Delta_{\max}\\ &=2\Delta_{\max}\langle W\rangle.\end{split}$ (26)

### v3

@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This paper is partially supported by Hong Kong Research Grants Council (RGC) grant C1042-23GF and grant #11205424 and Hong Kong Innovation and Technology Fund (ITF) grant MHP/061/23 and grant MHP/034/22.
Appendices Contents
@@ A.1 Neural Network Pruning
Neural network pruning [41, 28, 21] is a widely used technique for reducing the computational and memory footprint of neural networks by eliminating redundant parameters. Pruning methods can be categorized into structured pruning [36, 14, 32, 39], which removes entire structures like filters or layers, and unstructured pruning [12, 21], which targets individual weights. While structured pruning simplifies model deployment by maintaining regular architectures, it often suffers from significant performance degradation at higher sparsity levels (e.g., beyond 10% [39]), limiting its compression potential. In contrast, unstructured pruning can achieve sparsity levels exceeding 50% with minimal accuracy loss, making it a more flexible approach. For this reason, our discussion focuses primarily on unstructured pruning.
Traditional pruning pipelines follow a train-prune-retrain paradigm: a dense model is first trained, less important weights are pruned based on some criterion (e.g., magnitude [21, 12], gradients [41, 38], or Fisher information [52]), and the model is then fine-tuned to recover performance. However, this process is computationally expensive, as it requires training a dense model before pruning.
To avoid this overhead, recent work explores dynamic sparse training (DST) [37, 9, 11], where sparse topologies are optimized during training without ever densifying the model. While promising, existing DST methods rely on deterministic pruning-and-regrowth strategies, which can lead to high variance in model performance and inefficient exploration of sparse topologies, especially in federated learning scenarios with heterogeneous data and partial client participation. To mitigate these challenges, we propose TSAdj, a probabilistic adjustment mechanism that enables more stable and robust sparse topology evolution.
@@ A.2 Federated Pruning
Federated Pruning is a technique that applies neural network pruning within the Federated Learning (FL) process. To avoid the requirements of dense training in traditional pruning methods, some previous methods determine the sparse model topology before federated training and do not adjust it during the FL process, such as SNIP [29] and Synflow [56]. This approach of pruning at initialization may overlook crucial data information, potentially leading to suboptimal model structures.
Dynamic sparse training methods, such as those employed by PruneFL [24], adjust the model during FL rounds. However, these methods require the entire model gradient to be sent to the server to ensure final performance, resulting in high communication costs. FedDST [1] and FedSGC [57] alleviate this issue by applying dynamic sparse training on clients. FedTiny [18] and FedMef [19] require clients only to upload the Top gradients. Despite these advancements, these methods adjust the model topology based on limited information from the currently selected clients, leading to issues of greedy adjustment and unstable topology. FedRTS addresses these challenges by applying TSAdj, significantly mitigating the problems associated with limited client information and enhancing the robustness of model adjustments by leveraging probabilistic distributions.
@@ A.3 Combinatorial Multi-Arm Bandit
The Multi-Armed Bandit (MAB) problem [55] is a fundamental framework for decision-making under uncertainty, where an agent repeatedly selects among multiple actions ("arms") with initially unknown rewards to maximize cumulative gains. The Combinatorial Multi-Armed Bandit (CMAB) [5] extends this classic problem to scenarios where rewards depend on combinations of selected arms rather than individual choices.
Two primary approaches have emerged for solving CMAB problems: The first is the Combinatorial Upper Confidence Bound (CUCB) [5], which uses optimistic reward estimates to balance exploration and exploitation. The second is Combinatorial Thompson Sampling (CTS) [6, 59], which employs probabilistic sampling from posterior distributions to dynamically adjust its strategy.
While CUCB’s deterministic nature makes it theoretically appealing for analysis, its inherent optimism [5] can lead to persistently selecting clearly suboptimal arm combinations. This behavior prevents stable convergence, making CUCB less suitable in dynamic pruning where consistent performance is crucial, as will be discussed in detail in Sec. F.3.1. For this reason, we built our TSAdj method on the more adaptive CTS framework.
@@ B.1 TSAdj
In each loop $t$, the aggregated weights $W^{agg}_{t+1}$ will be calculated in each loop when the server receives the weights $W^{n}_{t+1}$ from each client $n\in C_{t}$, and TSAdj is invoked to adjust the sparse topology $m_{t}$ only in the outer loop. The outcomes $X_{t,i}$ for link $m_{i}$ are computed as described in Equation 3, which balances global and client-specific contributions. Specifically, outcomes $X^{agg}_{t,i}$ and $X^{n}_{t,i}$ are calculated based on Equations 4 on the server, using the aggregated weights $W_{t+1}$ and the uploaded weights $W^{n}_{t+1}$, while $X^{n}_{t,i}$ will be updated in the outer loop using the indices of the top $K-\kappa$ gradients $\mathcal{I}_{t+1}^{n}$ from each client $n$ based on Equation 5. Outcomes $X^{agg}_{t,i}$ and $X^{n}_{t,i}$ are combined to compute the final outcome $X_{t,i}$ for link $m_{i}$ in Equation 3, where $\gamma\in[0,1]$ controls the trade-off between global and client-specific contributions, and $p_{n}$ reflects the relative weight of client $n$ based on the size of its local dataset $D_{n}$. A larger $\gamma$ emphasizes client-specific observations, while a smaller $\gamma$ prioritizes global contributions.
The posterior probability distribution $P$ for each link $m_{i}$ in the topology $m$ is modeled as a beta distribution parameterized by $(\alpha_{i},\beta_{i})$. At initialization, $(\alpha_{i},\beta_{i})$ are set to small positive values (e.g., $\alpha_{i}=\beta_{i}=1$) to represent a uniform prior, ensuring that all links have an equal probability of activation at the outset. The beta parameters $(\alpha_{i},\beta_{i})$ for each link $m_{i}$ are then updated incorporating with the computed outcomes $X_{t,i}$, via Equation 6, where $\lambda$ controls the trade-off between exploration and exploitation by adjusting the impact of the current outcomes on the beta distribution’s update. A larger $\lambda$ increases the influence of the current loop’s outcomes $X_{t,i}$, accelerating the concentration of the Beta distribution towards a certain probability. Conversely, a smaller $\lambda$ preserves uncertainty to encourage exploration. Over time, as $\alpha_{i}$ and $\beta_{i}$ grow under the influence of $\lambda$ and $X_{t,i}$, the Beta distribution becomes sharper and more concentrated, reflecting an increasingly certain probability of selecting $m_{i}$ as an active link in the topology $m$.
@@ B.2 FedRTS
The training process is conducted over $T_{max}$ loops. At the beginning of each training loop $t$, the server employs a random sampling mechanism to select a subset of clients $C_{t}$ from the total $\mathcal{N}$ available clients for participation in the training process. This sampling process is conducted uniformly at random, ensuring that each client, indexed by $n$ from $1,\dots,\mathcal{N}$, has an equal probability of being chosen. This randomized partial participation mechanism not only avoids bias in client selection but also ensures scalability and reduces communication overhead, making it well-suited for large-scale distributed systems.
@@
Input: Initial model $W_{0}$, target sparsity $d^{\prime}$, learning rate $\eta$, maximum of training loops $T_{end}$, maximum of all loops $T_{max}$, adjustment interval $\Delta T$, training datasets $\{D_{n}\}_{n\in\mathcal{N}}$, the number of core links $\kappa$,
Output: Final sparse model $W_{T_{max}},m_{T_{max}}$
@@
for each loop $t=0,1,2,\dots,T_{max}-1$ do
  // The server does
@@
  Broadcast server sparse model ($W_{t}$, $m_{t}$) to all clients $n\in C_{t}$
@@
   // The client $n$ does
   $W_{t}^{n}$, $m_{t}^{n}$ $\leftarrow$ $W_{t}$, $m_{t}$ // Fetch global model
   $W^{n}_{t+1}\leftarrow$ $W_{t}^{n}-\eta\nabla f_{n}(W_{t}^{n},m_{t},D_{n})\odot m_{t}$
@@
    Sample batch data $\mathcal{B}_{n}\sim D_{n}$
    $G^{n}_{t+1}\leftarrow\nabla f_{n}(W_{t+1}^{n},m_{t},\mathcal{B}_{n})\odot m_{t}$
    $\mathcal{I}_{t+1}^{n}\leftarrow\{i|i\in\text{Top}(|G^{n}_{t+1}|,K-\kappa)\}$ // Top $K-\kappa$ magnitudes’ indices
@@
   $W^{agg}_{t+1},m_{t+1}\leftarrow\text{TSAdj}(\{W^{n}_{t+1}\}_{n\in C_{t}},\{\mathcal{I}_{t+1}^{n}\}_{n\in C_{t}}$) // TSAdj in Algorithm. 1.
@@
   // Update beta distribution based on weights’ outcomes
@@
   $m_{t+1}\leftarrow m_{t}$  // inner loop do not update topology
   $X^{agg}_{t},\{X^{n}_{t}\}_{n\in C_{t}}\leftarrow$ calculate with $W^{agg}_{t+1}$ and $\{W^{n}_{t+1}\}_{n\in C_{t}}$ by Eq. 4
   $X_{t}\leftarrow\gamma X^{agg}_{t}+(1-\gamma)\sum_{n\in C_{t+1}}p_{n}X^{n}_{t}$
   for each links $i=1,2,\dots,\langle X_{t}\rangle$ do
@@
After completing $T_{end}$ training loops, the server outputs the final globally aggregated weights $W_{T_{max}}$ and an optimized mask $m_{T_{max}}$, both derived from the iterative training involving all participating clients. The weights $W_{T_{max}}$ represent the learned parameters that effectively accommodate heterogeneous data distributions, while the topology $m_{T_{max}}$ preserves the robust and reliable structure of the model.
@@ C.1 Greedy Optimal Action
Theorem. (Upper Bound) Under assumptions 3.23.3 and 3.1 and with outcomes defined in Eq. 7, the regret $Reg(T)$ of TSAdj can be upper bounded by:
$\begin{split}Reg(T)&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2LK\epsilon)^{2}}+\left(2K+4\langle W\rangle+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}\\ &=O\left(\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{\Delta_{max}L^{2}\log T}{\Delta^{2}_{s,k}}\right).\end{split}$ (11)
for any $\epsilon$ such that $\forall s\not=S^{*}_{1}$ and $i\notin S^{*}_{k},\Delta_{s,k}>2LK\epsilon$, where $U$ is a universal constant.
Proof. For any arm $s\not=s^{*}_{1}$, we apply the exploration price $F(i)$ as
@@
Therefore, at the $t$-th iteration, there are two events
@@
Based on Eq. 13 and 14 , the regret $Reg(T)$ can be devided into three terms as
$\begin{split}Reg(T)&\leq\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{\neg A_{t}\land\neg B_{t}}\Delta_{S_{t}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{A_{t}}\Delta_{S_{t}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{B_{t}}\Delta_{S_{t}}\right].\end{split}$ (15)
We provide the Lemma 17,  23 and  26 to bound these three terms one by one. Therefore, we can obtain
$\begin{split}Reg(T)&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2LK\epsilon)^{2}}+\left(2K+4\langle W\rangle+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}\end{split}$ (16)
C.2.1 Lemmas
@@
$\begin{split}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{\neg A_{t}\land\neg B_{t}}\Delta_{S_{t}}\right]&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2LK\epsilon)^{2}}+\left(2K+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}\end{split}$ (17)
@@
[…]
@@
$\begin{split}\sum_{k=1}^{K}\mathbb{E}\left[\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{\neg A_{t}\land\neg B_{t}\land C_{t,k}}\right]&\leq\sum_{k=1}^{K}\mathbb{E}\left[\sum_{s\notin S^{*}_{k}}\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{s=s_{t,k}\land N_{t,s}\leq F(s)}\right]+\frac{UK}{\epsilon^{6}}\Delta_{max}\\ &\leq\mathbb{E}\left[\sum_{s\not=s^{*}_{1}}\sum_{t=1}^{T}\sum_{k=1}^{K}\Delta_{S_{t}}\cdot\mathbf{1}_{s=s_{t,k}\land N_{t,s}\leq F(s)}\right]+\frac{UK}{\epsilon^{6}}\Delta_{max}\\ &\leq\mathbb{E}\left[\sum_{s\not=s^{*}_{1}}\sum_{t=1}^{T}\Delta_{S_{t}}\cdot\mathbf{1}_{s\in S_{t}\land N_{t,s}\leq F(s)}\right]+\frac{UK}{\epsilon^{6}}\Delta_{max}\\ &\leq\sum_{s\not=s^{*}_{1}}F(s)\Delta_{s}^{\max}+\frac{UK}{\epsilon^{6}}\Delta_{max}\\ &\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2LK\epsilon)^{2}}+\frac{UK}{\epsilon^{6}}\Delta_{max},\end{split}$ (20)
@@
[…]
@@
$\begin{split}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{\neg A_{t}\land\neg B_{t}}\Delta_{S_{t}}\right]&\leq\sum_{s\not=s^{*}_{1}}\max_{k:s\notin S^{*}_{k}}\frac{6\Delta_{s}^{\max}L^{2}\log T}{(\Delta_{s,k}-2LK\epsilon)^{2}}+\left(2K+\frac{KU+8K\epsilon^{4}}{\epsilon^{6}}\right)\Delta_{max}.\end{split}$ (22)
@@
$\begin{split}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{A_{t}}\Delta_{S_{t}}\right]&\leq\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{A_{t}}\right]\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\mathbb{E}\left[\sum_{t=1}^{T}\mathbf{1}_{|\xi_{t,i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}}\right]\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\land|\xi_{t,i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\right)\Delta_{\max}\\ &=\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\cdot\mathbb{P}\left(|\xi_{t,i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\middle|N_{t,i}=\tau\right)\Delta_{\max}.\end{split}$ (24)
@@
$\begin{split}\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\cdot&\mathbb{P}\left(|\xi_{t,i}-\hat{\mu}_{t,i}|^{2}>\frac{3\log T}{2N_{t,i}}\middle|N_{t,i}=\tau\right)\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\sum_{\tau=1}^{T-1}\mathbb{P}\left(N_{t,i}=\tau\right)\cdot 2\exp(-3\log T)\Delta_{\max}\\ &\leq\sum_{i=1}^{\langle W\rangle}\sum_{t=1}^{T}\frac{2}{T}\Delta_{\max}\\ &=2\Delta_{\max}\langle W\rangle.\end{split}$ (25)